The Data of Macroe conomics- End of Chapter Problem A farmer grows a bushel of wheat and sells it to a miller for $1. The miller turns the wheat into flour and then sells the flour to a baker for $3. The baker uses the flour to make bread and sells the bread to an engineer for $6. When the engineer eats the bread, what is the value added by each person? What is the breads contribution to GDP? The farmers added value is S The millers added value is $ The bakers added value is $ The breads contribution to GDP is $

Need help, please show work
Answer #1

1) Value added by farmer = $1 (The value added will be equal to the amount at which it sold to the miller)

2) The miller’s added value = $(3-1) = $2 (The value added by the miller is the difference between the amount at which it bought wheat from the farmer and the amount at which it sold to the baker.)

3) The baker’s added value = $(6-3) = $3 (The value added by the baker is the difference between the amount at which it bought floor from the miller and the amount at which it sold bread to the engineer.)

4) The bread’ contribution to GDP = $(1+2+3) = $6 ( The contribution will be equal to the value added at each stage of production.)

* Please upvote.if satisfied.